From GlueXWiki
Jump to: navigation, search

June 28, 2012 FDC meeting


  1. FDC installation and cable routing (Tom Carstens)
  2. Production Construction Tracking (Dave)
    • Production status
    • Other: package transportation, subsystem installations
  3. Engineering update
    • Cooling system
  4. Electronics update (Chris)
  5. Cable assignment (Vlad)
  6. Gas system (Beni)
  7. Test set-up in 126 (Beni)
    • Cathode tests of the first package (Vlad)
  8. Other


Participants: Tom, Scott, Eugene, Tim, Slava, Mark, Dave, Chris, Nick, Vlad, Simon, Beni, and Lubomir. Many participants this time, some of them left after Tom's presentation.

FDC installation and cable routing

- Tom went through the slides linked above showing the FDC in three positions and on separate slides only the cable routing for these positions. The cables are divided in four quadrants and bundled together between the mesh cylinder (that separates the cables from the CDC) and the electronics racks. Left and right sides are symmetric. Two bundles run on the floor turn, below the platform, and go up through openings to the racks. The other two bundles will run over hangers attached to the racks and installation carts.

- The length of the bundles in the drawing is biggest for the initial position. While the FDC is being moved into the magnet, the top sections of the carts will be removed and the lower bundles will be lifted (there's enough slack) over the lower part of the carts. According to Tom this will be the most difficult part: the bundles have to be well supported to avoid deformation of the mesh cylinder. During this operation the bent position of the bundles will change all the time. Despite the stiffness of the bundle this will be possible (Tim, Tom). In the final position the length of the bundled section will be smallest, in reality the bundle length will not change, which means we have to find place for service loops. There will be additional loops of the individual cables either inside the bundles or close to the electronics.


- Dave: Installation of the fourth package was completed on Monday. Today Vlad noticed that one cathode in the middle was positioned incorrectly (15deg off). We discussed if it is possible to lift the top three cells together, correct the cathode position and then put the three cells back. For that we pull down the installation pins (one by one) to the right position so that separate only the top three cells and then insert from the top additional pins to keep to top part together. After the meeting this operation was done successfully. Three of the pins we did like this, for the other three we pushed the top pins further down so that it was easier to insert the top part in the right place. We can use this technique to separate part of the package if we have to fix something there.

- Dave showed the scheme (linked above) for the package refurbishment. As we did with the third and fourth packages we will install and test them in two halves. Techs working on the last two spare cathodes and preparing the last two spare wire frames for testing.

- Chris will submit PR for four PCB sets. We need to build (so far) one wire frame and for that we have a spare laminated G10 frame. If more wire frames are needed we will have to order more G10 frames, as well.

- Today the first package was moved back from EEL126 to Blue Crab for refurbishment. We are preparing for the installation of the subsystems on the third package. Will try to print the pre-amp labels from file and put them on the cards before the installation.

Engineering update

- Eugene asked us before if we can use water for cooling instead of Fluorinert and if there are examples for that. At ATLAS all the cooling systems that are not serviceable use some type of Fluorinert, while for the outer system water cooling is used as well.

- We discussed what will happen in case of water leakage and if there's a way to put containers below the manifolds. The main question, if it is safe to use water, remains. Certainly this has to be discussed with Bill and Fernando.

Electronics update

- Chris: delay with the grounding clips, they were supposed to be done by today, but the company answered they will send us first article sometimes next week.

Cable assignment

- Vlad: final scheme linked above. Same logic for distributing the loads within a package, but this time the packages are separated, i.e no mixture of cables from different packages.

Gas system

- Beni: Narciso has a preliminary scheme and scheduled a meeting beginning of next month.

Test set-up at 126

- Beni: working on CODA-Light for the two-crate system, will move to Linux ROCs. Eugene: why not to use real CODA?

- Vlad tested the cathodes of five cells (the 55Fe source is not strong enough for the sixth cell). All channels working, but same complains as before: need to fix the cooling tubes somehow to the chamber. We will attach them to the holes at the perimeter of the frames with cable ties. Problems with the pins of two LV connectors.